OpenStudy (anonymous):

Determine the perimeter of ABCD. (Points : 2) 5 28 18 10

OpenStudy (anonymous):

@jim_thompson5910 i need help with this one its hard

jimthompson5910 (jim_thompson5910):

what's the length of AB

OpenStudy (anonymous):

2

jimthompson5910 (jim_thompson5910):

going from -5 to 4 on the number line, you are walking _____ steps

OpenStudy (anonymous):

10 steps

jimthompson5910 (jim_thompson5910):

close, 9 steps actually 4 - (-5) = 4+5 = 9

jimthompson5910 (jim_thompson5910):

what's the length of AD

OpenStudy (anonymous):

5

jimthompson5910 (jim_thompson5910):

add these two lengths together 9+5 = 14 then double this sum 14*2 = 28 that's your perimeter

jimthompson5910 (jim_thompson5910):

alternatively, you can just add up all the sides or you can double the sides first, then add

OpenStudy (anonymous):

cool thanks :)